OpenStudy (anonymous):

WILL MEDA! I think I got it but idk. plz help me What is the value of a? 9a-3(a-6)=-6 Is It -24?

OpenStudy (anonymous):

@goldendragon01 ?

OpenStudy (anonymous):

9a - 3(a - 6) = - 6 9a -3a + 18 = -6 6a + 18 = - 6 -18 -18 6a = -24 -------- 6 6 a = - 4

OpenStudy (anonymous):

I failed it anyways -_-

OpenStudy (anonymous):

Was my answer wrong?

OpenStudy (anonymous):

yes. surprisingly. :/ I sear school is rigged lol @HeatDude

OpenStudy (anonymous):

lol that is weird